# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v1

@@ B.3 Analysis of Different Score Functions
A simple case study.
@@
Formula to compute MGFs.
@@
  • Suppose that we already show the statement holds for $m\leq\lfloor\frac{1}{1-\Delta}\rfloor$, we now to show it also holds for $m-1$. For notation simplicity, we let ${\mathcal{P}}_{m}\subset{\mathcal{P}}_{\Delta}$ denote the set of NTP distributions which having exact $m$ coordinates equal to $1-\Delta$. Let $\bm{P}\in{\mathcal{P}}_{m-1}$ be the NTP distribution under consideration. Given its coordinates are decreasingly ranked, we must have $1-\Delta>P_{m}\geq P_{m+1}\geq\cdots\geq P_{|{{\mathcal{W}}}|}$.
@@
B.5 Proof of Theorem 3.3
@@
The constant $10^{-3}$ used in the above proof is not essential. For a more accurate characterization, we use the python SciPy package to solve the root of $R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ars}})=R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{log}})$, the numerical result suggests that there is only one root with value $0.17756080525215662$ when $\Delta\geq 0.001$.
B.6 Analysis of Efficiency Gap
@@
B.7 ${\mathcal{P}}_{\Delta}$-Efficiency for the Baby Watermark
@@
$R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{id}})$ $=-\inf_{\theta\geq 0}\log\biggl{[}\frac{1}{\theta}\biggl{\{}\frac{{\mathrm{e}}^{\theta(1-2\Delta)}+{\mathrm{e}}^{-\theta(1-2\Delta)}}{2}-{\mathrm{e}}^{-\theta}\biggr{\}}\biggr{]}.\qed$
@@ C.1 Distribution Characterization
$=\frac{1}{|{{\mathcal{W}}}|!}\sum_{\pi\in\Pi}{\mathbb{P}}\left(U_{t}\in\left(a_{\pi,{w}-1},a_{\pi,{w}}\right]\cap[0,r]\right).\qed$
@@ C.2 Key Lemma for Asymptotic Analysis
$\mathcal{Q}_{\Delta}=\left\{\bm{P}:P_{(1)}=1-\Delta,\log|{{\mathcal{W}}}|\cdot P_{(2)}\leq\varepsilon_{|{{\mathcal{W}}}|}\right\}.$ (38)
@@
$=:{\mathcal{T}}_{1}+{\mathcal{T}}_{2}+{\mathcal{T}}_{3}.$ (40)
In the following, we analyze the three terms ${\mathcal{T}}_{1},{\mathcal{T}}_{2}$ and ${\mathcal{T}}_{3}$ in (40) respectively.
For the term ${\mathcal{T}}_{1}$.
@@
$\qquad\leq O(1)\cdot\left(\frac{1}{|{{\mathcal{W}}}|}+\sqrt{P_{2}\log|{{\mathcal{W}}}|}\right),$ (43)
@@
xholds uniformly over $\bm{P}\in\mathcal{Q}_{\Delta}$ and $\Delta\in[0,1-\frac{1}{|{{\mathcal{W}}}|}]$ in the sense that
@@
Visualization of weak convergence on an extreme case.
@@ D.1 Simulation Setup
Choice of pseudorandom variable.
@@
Computation of critical values.
@@
Additional results for other choices of $b$.
@@ D.2 Real-world Experimental Details
Extended results for the Sheared-LLaMA-2.7B model.

### v2

@@ B.3 Analysis of Different Score Functions
A simple case study
@@
Formula to compute MGFs
@@
  • Suppose that we already show the statement holds for $m\leq\lfloor\frac{1}{1-\Delta}\rfloor$, we now need to show that it also holds for $m-1$. For notation simplicity, we let ${\mathcal{P}}_{m}\subset{\mathcal{P}}_{\Delta}$ denote the set of NTP distributions which having exact $m$ coordinates equal to $1-\Delta$. Let $\bm{P}\in{\mathcal{P}}_{m-1}$ be the NTP distribution under consideration. Given its coordinates are decreasingly ranked, we must have $1-\Delta>P_{m}\geq P_{m+1}\geq\cdots\geq P_{|{{\mathcal{W}}}|}$.
@@
B.5 Results for a General Probability Set
In the following, we consider a new, general distribution class ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$. For two real numbers $10\leq\Delta_{2}<\Delta_{1}<$ 1, we define the following probability class:
${\mathcal{P}}_{\Delta_{1},\Delta_{2}}=\left\{\bm{P}\in{\mathrm{Simp}}({{\mathcal{W}}}):P_{(1)}\leq 1-\Delta_{1},P_{(2)}\leq 1-\Delta_{2}\right\},$ (34)
where $P_{(i)}$ is the $i$-th largest probability in the NTP distribution $\bm{P}=(P_{1},\cdots,P_{|{{\mathcal{W}}}|})$. Notably, ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ is invariant under permutation, meaning that for any permutation $\pi$, if $\bm{P}\in{\mathcal{P}}_{\Delta_{1},\Delta_{2}}$, then $\pi(\bm{P}):=(P_{\pi(1)},\ldots,P_{\pi(|{{\mathcal{W}}}|)})$ also belongs to ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$. Furthermore, ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ is a convex polytope, defined as a compact convex set with a finite number of extreme points. The permutation invariance allows us to precisely capture its set of extreme points.
Specifically, we establish the following lemma:
Lemma B.3.
For $10\leq\Delta_{2}<\Delta_{1}<$ 1, the set of extreme points of ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ is
$\mathrm{Ext}({\mathcal{P}}_{\Delta_{1},\Delta_{2}})=\Pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star})$
where $\Pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}):=\left\{\pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}):\pi~{}\text{is a permutation on $\{1,2,\ldots,|{{\mathcal{W}}}|\}$}\right\}$ and $\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}$ is the least-favorable NTP distribution in ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ defined by
$\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}=\biggl{(}1-\Delta_{1},\underbrace{1-\Delta_{2},\ldots,1-\Delta_{2}}_{\lfloor\frac{\Delta_{1}}{1-\Delta_{2}}\rfloor~{}\text{times}},\Delta_{1}-(1-\Delta_{2})\cdot\left\lfloor\frac{\Delta_{1}}{1-\Delta_{2}}\right\rfloor,0,\ldots\biggl{)}.$
Following a similar line of reasoning as in our main analysis, we could show that
$h_{\mathrm{gum},\Delta_{1},\Delta_{2}}(r)=\log\frac{{\mathrm{d}}\mu_{1,\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}}}{{\mathrm{d}}\mu_{0}}(r)$
is the optimal score function for the Gumbel-max watermark when using ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ as the prior probability set.
However, this probability set is less practical because it requires specifying two real numbers $\Delta_{1},\Delta_{2}$. Therefore, we include this discussion as an extension, illustrating how our analysis and method can be extended to other permutation-invariant sets ${\mathcal{P}}$. In the main text, we focus on the $\Delta$-regular prior set ${\mathcal{P}}_{\Delta}$, which represents a special case of ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$. This case occurs when we set $\Delta_{1}=\Delta$ and $\Delta_{2}=0$, making it perhaps the simplest form of this class.
Proof of Lemma B.3.
The proof strategy closely follows the approach used in Lemma 3.4, with a few key distinctions that we detail here.
First, it is straightforward to see that $\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}$ and its permutations are all extreme points since they reach the boundary conditions $P_{(1)}\leq 1-\Delta_{1}$ and $P_{(2)}\leq 1-\Delta_{2}$. This result can be easily verified by using the definition of extreme points so we omit them for simplicity. Consequently, we have $\Pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star})\subseteq\mathrm{Ext}({\mathcal{P}}_{\Delta_{1},\Delta_{2}})$.
To prove the reverse inclusion, $\mathrm{Ext}({\mathcal{P}}_{\Delta_{1},\Delta_{2}})\subseteq\Pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star})$, it suffices to demonstrate that any point in ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ lies within the convex hull of $\Pi(\bm{P}_{\Delta_{1},\Delta_{2}}^{\star})$. This can be shown by applying the following lemma. By the terminology introduced in Lemma B.4, we know that $\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}$ majorizes any point in ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$, implying that any point in ${\mathcal{P}}_{\Delta_{1},\Delta_{2}}$ can be expressed as a convex combination of $\bm{P}_{\Delta_{1},\Delta_{2}}^{\star}$ and its permutations. Thus, the proof is complete.
Lemma B.4 (Lemma 2.2 in [34]).
Given two vectors ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{d}$, we say ${\bm{x}}$ majorizes ${\bm{y}}$, a property we denote by ${\bm{x}}\geq_{m}{\bm{y}}$, if $\sum_{i=1}^{j}x_{(i)}\geq\sum_{i=1}^{j}y_{(i)}$ for all $j=1,\ldots,d$ and $\sum_{i=1}^{d}x_{(i)}=\sum_{i=1}^{d}y_{(i)}$. Here $x_{(i)}$ is the $i$-th largest component of ${\bm{x}}$. If ${\bm{x}}\geq_{m}{\bm{y}}$, then ${\bm{y}}$ is a convex combination of ${\bm{x}}$ and its permutations.
B.6 Proof of Theorem 3.3
@@ B.5 Results for a General Probability Set
The constant $10^{-3}$ used in the above proof is not essential. For a more accurate characterization, we use the python SciPy package to solve the root of $R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{ars}})=R_{{\mathcal{P}}_{\Delta}}(h_{\mathrm{log}})$, the numerical result suggests that there is only one root whose value is $0.17756080525215662$ when $\Delta\geq 0.001$.
B.7 Analysis of Efficiency Gap
@@
B.8 ${\mathcal{P}}_{\Delta}$-Efficiency for the Baby Watermark
@@ C.1 Distribution Characterization
$=\frac{1}{|{{\mathcal{W}}}|!}\sum_{\pi\in\Pi}{\mathbb{P}}\left(U_{t}\in\left(a_{\pi,{w}-1},a_{\pi,{w}}\right]\cap[0,r]\right).$
@@ C.2 Key Lemma for Asymptotic Analysis
$\mathcal{Q}_{\Delta}=\left\{\bm{P}:P_{(1)}=1-\Delta,\log|{{\mathcal{W}}}|\cdot P_{(2)}\leq\varepsilon_{|{{\mathcal{W}}}|}\right\}.$
@@
$=:{\mathcal{T}}_{1}+{\mathcal{T}}_{2}+{\mathcal{T}}_{3}.$
In the following, we analyze the three terms ${\mathcal{T}}_{1},{\mathcal{T}}_{2}$ and ${\mathcal{T}}_{3}$ respectively.
For the term ${\mathcal{T}}_{1}$
@@
$\qquad\leq O(1)\cdot\left(\frac{1}{|{{\mathcal{W}}}|}+\sqrt{P_{(2)}\log|{{\mathcal{W}}}|}\right),$ (42)
@@
Visualization of weak convergence on an extreme case
@@ D.1 Simulation Setup
Choice of pseudorandom variable
@@
Computation of critical values
@@
This estimate is consistent in the sense that we have $\lim_{n\to\infty}\gamma_{n,\alpha}/\hat{\gamma}_{n,\alpha}\to 1$ for any $\alpha\in(0,1)$ and any score function satisfying $\operatorname{\mathbb{E}}_{0}|h(Y)|^{2}<\infty$.
Remark D.1.
Hash collisions occur when the computed pseudorandom numbers are not truly i.i.d. It makes the way using $\hat{\gamma}_{n,\alpha}$ to control Type I error become invalid. This issue is well illustrated in Figure 1 of [22], where it is shown that the empirical false positive rates for common watermarks often exceed theoretical predictions. The primary reason for this discrepancy is that their pseudorandom numbers rely only on local information, i.e., $\zeta_{t}={\mathcal{A}}(w_{(t-m):(t-1)},{\mathtt{Key}})$ in our notation. Consequently, if there exists any $t\neq t^{\prime}$ such that $w_{(t-m):(t-1)}=w_{(t^{\prime}-m):(t^{\prime}-1)}$, then $\zeta_{t}=\zeta_{t^{\prime}}$, leading to the same pseudorandom numbers appearing multiple times. This violates the assumption of true i.i.d. nature and thus our working hypothesis. To mitigate this, [22] proposed heuristic methods to address such repetitions (see their Part C in Section III). By implementing these methods, they were able to improve Type I error control, as demonstrated in the rightmost panel of Figure 1.
However, in our paper, we did not consider the issue of hash collisions, as this issue is orthogonal to the main results of our paper, and addressing it is beyond the scope of our current work. Instead, our approach assumes $\zeta_{t}={\mathcal{A}}(w_{1:(t-1)},{\mathtt{Key}})$, which effectively corresponds to the case where $m=\infty$. Given that the hash function ${\mathcal{A}}$ is highly sensitive to its inputs, even slight differences in these inputs result in outputs that are statistically independent. Therefore, as long as $t\neq t^{\prime}$, we have $w_{1:(t-1)}\neq w_{1:(t^{\prime}-1)}$, ensuring that $\zeta_{t}\perp\zeta_{t^{\prime}}$. This implies that our pseudorandom numbers are truly i.i.d., allowing us to confidently use $\hat{\gamma}_{n,\alpha}$ to control the Type I error.
Additional results for other choices of $b$
@@ D.2 Real-world Experimental Details
Extended results for the Sheared-LLaMA-2.7B model
@@ D.3 Details of Figure 1
D.4 Discussion on the Selection of $\Delta$
In practice, we choose $\Delta$ using prior knowledge. It is important to note that there is a trade-off when selecting the value of $\Delta$.
Assume that an exact $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ belongs to ${\mathcal{P}}_{\Delta}$, and we lack information about the remaining NTP distributions. Consequently, the rest are considered to be in the full class ${\mathcal{P}}_{0}={\mathrm{Simp}}({{\mathcal{W}}})$, which includes all categorical distributions over the vocabulary ${{\mathcal{W}}}$. Proposition 2.1 indicates that $\lim\sup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1})=0)^{1/n}\leq\mathrm{e}^{-\gamma R_{{\mathcal{P}}_{\Delta}}(h)}$. This inequality is tight in the worst-case scenario, implying that without additional knowledge, this is the best convergence rate we can expect.
Theoretically, we view $\gamma$ as a function of $\Delta$, interpreting it as the fraction of $\Delta$-regular NTP distributions among $\bm{P}_{1},\ldots,\bm{P}_{n}$. Clearly, $\gamma(\Delta)$ is a decreasing function of $\Delta$. As $\Delta$ increases, we impose a stricter condition on the largest probability in $\bm{P}$, requiring $\max_{{w}\in{{\mathcal{W}}}}P_{{w}}\leq 1-\Delta$, resulting in a smaller fraction satisfying this condition. However, as shown in Figure 4 (or the middle panel of Figure 12), $R_{{\mathcal{P}}_{\Delta}}(h)$ is an increasing function of $\Delta$. Therefore, the final efficiency rate, $\gamma(\Delta)R_{{\mathcal{P}}_{\Delta}}(h)$, is a complex function of $\Delta$, neither strictly increasing nor decreasing. This implies a trade-off in choosing $\Delta$, and an optimal choice exists.
Figure 12: Illustration of $\gamma(\Delta)$ (left), $R_{{\mathcal{P}}_{\Delta}}(h)$ (middle), and $\gamma(\Delta)\cdot R_{{\mathcal{P}}_{\Delta}}(h)$ (right) as functions of $\Delta$. Since $\Delta$ is typically small in practice, we focus on the range $\Delta\in[0.001,0.5]$. The experimental setup is consistent with that of Figure 1.
We support the above analysis with further empirical study, following the same setup as in Figure 1 and presenting the results in Figure 12. Its left panel shows the empirical estimation of $\gamma(\Delta)$, which is clearly a decreasing function of $\Delta$. The middle panel, essentially a zoomed-in version of Figure 4, is presented here within the interval $\Delta\in[0.001,0.5]$ for reader convenience. The $\arg\max$ denotes the value of $\Delta$ at which the value of the $y$-axis reaches its maximum. Since theoretically $R_{{\mathcal{P}}_{\Delta}}(h)$ is an increasing function of $\Delta$, we find that $\arg\max\equiv 0.5$ for all score functions. However, when we focus on the product $\gamma(\Delta)\cdot R_{{\mathcal{P}}_{\Delta}}(h)$, it typically increases at first and then decreases. This can be validated by observing the value of $\arg\max$. In the right panel of Figure 12, the computed values of $\arg\max$ for all score functions shift from $0.5$, indicating that the inclusion of $\gamma(\Delta)$ indeed impacts the final efficiency. We highlight that in our particular example, the effect of $\gamma(\Delta)$ seems limited, as the shift in $\arg\max$ is moderate. However, there are cases where this deviation can be significant. Hence, choosing a good value of $\Delta$ would be a case-by-case argument, but the idea of choosing an optimal $\Delta$ would always be applicable.